Find the measure of an angle if its measure is four times the measure of its complement?: Find the measure of an angle if its measure is four times the measure of its complement?
Answer by jojo14344(814)

We know complement angles form 90 degrees
Let's make 2 angles as angleA & it's complement as angleB
So, (angleA)+(angleB)=90 -----------> working eqn
but, angleA=4*angleB,= A=4B, then
4B+B=90 -----------> cross(5)B/cross(5)=cross(90)18/cross(5)
B=18deg
And, A=4*18=72deg
In doubt? go back working eqn,
72+18=90
90=90
Thank you,
Jojo